OpenStudy (anonymous):

Find the 6th term of the sequence an=an-1 +8 where a1=10 please show work

OpenStudy (tkhunny):

Well, you had better start showing your work! a2 = a1 + 8 a1 = 10 a2 = 10 + 8 = 18 Okay, you do the next 4 and you'll be done.

OpenStudy (anonymous):

why did you put a2

OpenStudy (anonymous):

see n in this case denotes the number of terms....

OpenStudy (anonymous):

so for finding 6th term ....put n=6 a6=a5+8

OpenStudy (anonymous):

oh okay so now what

OpenStudy (tkhunny):

It's a sequence. That's what the notation means. It should be a subscript. \(a_{2} = f(a_{1}) = a_{1} + 8\) The formula is this: \(a_{n} = f(a_{n-1}) = a_{n-1} + 8\) This says only that the value of each member of the sequence is eight (8) greater than the previous member.

OpenStudy (anonymous):

what now

OpenStudy (tkhunny):

Whaat now, what? Do you have \(a_{2}\)? If so, calculate \(a_{3}\). Do you have \(a_{3}\)? If so, calculate \(a_{4}\) and so on...